Verify: csc⁴x-2csc²x+1=cot⁴x
1
Expert's answer
2019-12-16T09:23:27-0500


We need to verify "csc^4 x - 2 csc^2 x + 1 = cot^4 x"


We know "a^2 - 2a + 1 = (a - 1)^2" and


"a^4 - 2 a^2 + 1 = (a^2 -1)^2"

"csc^2 x - 1 = cot^2 x"

Now,

apply the above formula for this Question



"csc^4 x - 2 csc^2 x + 1 = (csc^2 - 1) ^2"


"=(cot^2 x)^2 = cot^4 x"

Answer: Yes, it was proved that



"csc^4 x - 2 csc^2 x + 1 = cot^4 x"